OpenStudy (anonymous):

What's the percent of change for 18. To 10

hero (hero):

I gave you the formula for this.

OpenStudy (anonymous):

Is it 80%

hero (hero):

Did you input the values into the formula?

OpenStudy (anonymous):

I think so is it 80

hero (hero):

I'm going to go over this one last time

hero (hero):

did you delete the previous question?

OpenStudy (slaaibak):

\[{{new - old} \over old} * 100\]

hero (hero):

The general formula for percent change is as follows: \[\% \Delta = \frac{|a-b|}{a} \times100\] where a = original amount, b = new amount In this case, a = 18 , b = 10 Plug the numbers in and solve.

OpenStudy (anonymous):

Ok what do you come up with

hero (hero):

you first. I did the last one for you already.

hero (hero):

This involves simple arithmetic

OpenStudy (anonymous):

I said it already if it's not 80 then is it 44.4 repeating

hero (hero):

Yes, it is 44.4%

hero (hero):

It's not 80. Never was, never will be. How you got 80 perplexes me.

hero (hero):

No, it stupifies me.

OpenStudy (anonymous):

Ok that's what I though but then the first one you didn't make sense

hero (hero):

I posted the same exact formula. Nothing changed.

hero (hero):

Don't say what I wrote doesn't make sense. It's not me that isn't making sense.

OpenStudy (anonymous):

So the first answer was 33.3 repeating or 25 as you said

hero (hero):

Just 25% not 33 Do you understand what the absolute value symbols mean?

hero (hero):

|| means |-x| = x In other words |-5| = 5 So basically, when you evaluate using the absolute value, you drop the negative and use only the positive value. How you're getting two values...I don't know. Don't pay attention to that other guy's formula. It will only confuse you.

OpenStudy (anonymous):

It was 15 to 20 20-15 is 5 5*100 500/15 is 33.3

hero (hero):

5/20 = 1/4 = 25% That's the solution to the first one. You're doing some incorrect math there. In the first one, a = 15, b = 20 a = the original amount, b = new amount 15 is the original amount, 20 is the new amount Don't confuse yourself.

hero (hero):

You multiply by 100 LAST

hero (hero):

You have to divide before multiplying by 100

OpenStudy (anonymous):

Ok I'll just go with your answer and when we check it in class tomorrow I'll just see

hero (hero):

Yes, and make sure you get back to me when you find that 25% is the correct answer.

OpenStudy (anonymous):

Oh yea I'll get back to you

hero (hero):

Hmm, I just checked my own answer. It really is 33.3%

hero (hero):

I take my foot out of my mouth. You are correct

hero (hero):

Nevertheless, the formula is correct. Use it.

OpenStudy (anonymous):

Oh I've been using it and I have been getting the correct answers

hero (hero):

Great. That's all that matters.

OpenStudy (anonymous):

Btw this made my day it was fun arguing with you thanks

hero (hero):

lol

hero (hero):

Yeah, I'm not perfect

OpenStudy (anonymous):

Oh yea I know

OpenStudy (slaaibak):

Hero! What rubbish are you posting here? If |-x| = -x IFF x<= 0 Otherwise, |-x| = x if x>=0

OpenStudy (slaaibak):

I'd rather define it as |-x| = -x if x<0 and |-x| = x if x>=0